OpenStudy (tanyatazf):

Write each exponential equation in its equivalent logarithmic form. 15=radical 225

OpenStudy (anonymous):

You can write it as: \[\large 15 = (225)^{\frac{1}{2}}\] Now use the Logarithmic Formula: \[\large \color{green}{Log_ab = x \implies a^x = b}\]

OpenStudy (anonymous):

So, reverse will be: \[\large \color{green}{a^x = b \implies Log_ab = x}\]

OpenStudy (anonymous):

So, can you do it now??

OpenStudy (tanyatazf):

so log 225 15=1/2?

OpenStudy (anonymous):

Yes you have done it very correctly..

OpenStudy (tanyatazf):

thank you!

OpenStudy (anonymous):

\[\large Log_{225}15 = \frac{1}{2}\]

OpenStudy (anonymous):

And last but not the least: \[ \huge \color{green}{\textbf{Welcome To Openstudy..}} \color{blue}{{\checkmark}}\]